2zx2y=18 where x=1,z=y
Answer by RAY100(1637)

2zx2y=18,,,but z=y,,,x=1,
.
2(y)(1)2y=18
.
4y^2 =18
.
y^2 =18/4 =4.5
.
y= +/-2.12
.